OpenStudy (anonymous):

What is the average of 52?

OpenStudy (anonymous):

I think you left out something... the "average of 52" doesn't really make any sense

OpenStudy (anonymous):

The whole question is: Seven more than the average of 4 numbers is 52. Find the sum of the four numbers.

OpenStudy (anonymous):

ok, that makes more sense. write the sentence as math: 7 + (average of 4 numbers) = 52 Then you can simplify and find the average of 4 numbers. Knowing the average of 4 numbers, can you figure out the sum? Play around with an easy example... the average of two numbers like 10 and 20... to get the hang of it.

OpenStudy (anonymous):

Ok I'm helping out my brother on his homework and Im not sure how to explain this to him.

OpenStudy (anonymous):

The point is, you don't have to know what the 4 numbers are... any four numbers that average 52 will add up to 4 times 52 or 208, because when you divide that sum by 4, you get 52.

OpenStudy (anonymous):

i don't need the answer i just need the equation for the answer

OpenStudy (anonymous):

Average = sum/(number of terms) so it's: 7 + (sum/4) = 52 sum/4 = 52-7 sum/4 = 45 sum = 45 * 4 = 180

OpenStudy (anonymous):

Also, sorry for earlier confusion.... my example of 208 was based on average of 52, but as you see in the equation above, it should be the average is 45, because the problem says "7 more than the average is 52"

OpenStudy (anonymous):

ok thanks